# Energy Window Augmented Plane Waves Approach to Density Functional Theory

### v5

@@ Abstract
In this work we present a new method for basis set generation for electronic structure calculations of crystalline solids. This procedure is aimed at applications to Density Functional Theory (DFT). In this construction, Energy Window Augmented Plane Waves (EWAPW), we take advantage of the fact that most DFT calculations use a convergence loop in order to obtain the self consistent eigenstates of the final (converged) Kohn Sham (KS) Hamiltonian. Here we propose that, for the basis used at each step of the self consistency iteration, we use the previous eigenstate basis, in the interstitial region, and augment it, inside each Muffin Tin (MT) sphere, with the solution to the spherically averaged KS Hamiltonian for the linearization energy of the energy window which contains the energy of that previous eigenstate. Indeed, to reduce the number of times the spherically averaged KS potential needs to be solved inside the MT spheres it is advantageous break up the spectrum into non-overlapping intervals, windows, and solve the spherically averaged KS Hamiltonian inside the MT region only once per window per angular momentum channel (at the linearization energy relevant to that window, usually near the middle of the window). For practical applications it is reasonable to have on the order of ten to one hundred windows. At each step of the iteration of the solution of the KS equations the EWAPW basis is that of near eigenstates of the KS Hamiltonian for that iteration.
@@ I Introduction
The usefulness of a basis set for all electron calculations of crystalline solids within Density Functional Theory (DFT) methods is mainly determined by 1) its linearization error (that is energy errors for eigenstates of the Kohn Sham (KS) Hamiltonian inside the valence band), 2) its ability to handle semi-core states - that is the situation when there is more then one relevant band (not just the valence band) just at or below the Fermi energy, 3) the effectiveness with which the basis set converges to the energy obtained in the infinite basis limit as a function of the number of basis elements for the solution of the KS equations, $N_{tot}$ or how hard or soft the basis is, 4) the efficiently with which one can write down the KS Hamiltonian in that basis, amongst other things [1, 2, 3]. There are now several known competitors, in the literature, for which basis set is optimal. These can be divided into mostly plane wave basis methods [1, 2, 8, 9, 5, 6, 3, 4, 7] and mostly Muffin Tin (MT) basis sets [2, 11, 4, 12, 13, 15, 10, 14]. In this work we focus on the plane wave methods [1, 2, 8, 9, 5, 6, 3, 4, 7]. Nominally (without augmentation or pseudization), within the plane wave method, the basis set is very simple but is extremely inefficient at obtaining good eigenstates of the KS Hamiltonian for reasonably small wave vector cutoffs - is extremely hard. Indeed near the nuclei, where the electron density is high and highly concentrated (oscillating), the KS potential has high wave vector components so that the eigenstates of the KS Hamiltonian have high wave vector components as well. This means that impractically large numbers of plane waves are needed to accurately describe the environment near atomic nuclei, making the situation generically numerically impossible. There are two main approaches to cure this problem - that is still use mainly plane wave basis sets: pseudopotentials [16, 17, 2, 3] and augmentation methods [4, 2, 18, 8, 9, 19, 20, 21, 7]. In this work we focus on augmentation methods.
@@
One solution to this glueing problem is the Linearized Augmented Plane Waves (LAPW) basis set [4, 2, 1]. Here plane waves are augmented within the MT sphere by a linear combination of the solutions to the spherically averaged KS Hamiltonian (see Eq. (1)) at the linearization energy $E_{L}^{n}$ (typically the middle of the valence band) $u_{l\alpha}^{n}\left(r,E_{L}^{n}\right)$ and its derivative with respect to energy $\dot{u}_{l\alpha}^{n}\left(r,E_{L}^{n}\right)=\frac{\partial}{\partial E}u_{l\alpha}^{n}\left(r,E_{L}^{n}\right)$. This is done for all angular momentum channels $lm$ and all atoms $\alpha$. Whereby, by matching the MT wavefunctions to the planewave wavefunction, that is demanding the total function be continuous and continuously differentiable - at the MT sphere radius [4, 2, 1, 5, 6] - we obtain a efficient smooth basis set. The LAPW basis set may be written as:
@@
Where $IR$ is the interstitial, $MT_{\alpha}$ is the $\alpha$th MT sphere, $\Omega$ is the volume of the unit cell, $\mathbf{k}$ is a point in the first Brillouin, $\mathbf{G}$ is a reciprocal lattice vector, $Y_{lm}$ are spherical harmonics, $a_{\mathbf{k}\mathbf{G}}^{lm\alpha n}$ and $b_{\mathbf{k}\mathbf{G}}^{lm\alpha n}$ are the matching coefficients that insure the wavefunction is continuous and continuously differentiable. Furthermore $a_{\mathbf{k}\mathbf{G}}^{lm\alpha n}$ and $b_{\mathbf{k}\mathbf{G}}^{lm\alpha n}$ can be computed using the relation [22]:
which decouples the matching problem into separate angular momentum channels. Here $j_{l}$ are the spherical Bessel functions. There is a cutoff $l_{max}\sim 8-12$ which comes about because of the negligibly small values of the spherical Bessel functions, for large $l$, inside the MT spheres. This allows for efficient description of the valence band which is adjustable to the form of the KS potential near the nuclei. This idea can be extended to Lattice Adapted Linearized Plane Waves ($\left(LA\right)^{2}PW$) which have the form [18]:
@@
for $\nu=j$. One then truncates the basis set with number of basis functions $j\ll N_{\mathbf{G}}$, as one has a basis of near eigenvectors in the IR, and runs the KS loop to convergence. This greatly reduces the basis set size needed for the iterative calculations after the first iteration [18]. These methods can be further improved as in Quadratically Augmented Plane Waves QAPW [20, 21] where $\ddot{u}_{l\mu}^{n}\left(r,E_{L}^{n}\right)=\frac{\partial^{2}}{\partial E^{2}}u_{l\mu}^{n}\left(r,E_{L}^{n}\right)$ is also used inside the MT sphere and one more derivative (with respect to the radial co-ordinate) is matched to the plane waves in the IR which reduces linearization errors. However, the more derivatives are matched the more the radial wavefunction inside the MT sphere look like Bessel functions (indeed in the limit where an infinite number of derivatives is matched the radial wavefunction must be exactly Bessel) the higher cutoff, the harder the basis becomes [7, 18]. Furthermore if there are several relevant bands one can use Local Orbitals (LO or lo) basis wave functions [8, 1, 9, 23, 18, 7]. Many LO like basis wavefunctions are of the form:
$\phi_{LO}^{nl\alpha}=\left\{\begin{array}[]{cc}0&\mathbf{r}\in IR\\ \left[\bar{a}^{l\alpha n}u_{l\alpha}^{n}\left(r,E_{L}^{n}\right)+\bar{b}^{l\alpha n}\dot{u}_{l\alpha}^{n}\left(r,E_{L}^{n}\right)+\bar{c}^{l\alpha n}\tilde{u}_{l\alpha}^{n}\left(r\right)\right]Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\alpha}}\right)&\mathbf{r}\in MT_{\alpha}\end{array}\right.$ (7)
Where the wavefunction is chosen to be continuous and continuously differentiable using the coefficients $\bar{a}^{l\alpha n}$, $\bar{b}^{l\alpha n}$ and $\bar{c}^{l\alpha n}$. There are many choices for $\tilde{u}_{l\alpha}^{n}$. For Higher Derivative Local Orbitals [23, 25, 7, 18, 24] (HDLO) $\tilde{u}_{l\alpha}^{n}=\ddot{u}_{l\alpha}^{n}\left(r,E_{L}^{n}\right)$, while for Higher Energy Local Orbitals [23, 25, 7, 18, 24] (HELO) the $\tilde{u}_{l\alpha}^{n}$ is an eigenstate of the spherically averaged KS Hamiltonian (inside $MT_{\alpha}$) with energy $E_{HELO}^{nl}$ such that
@@
In this work we will not be proposing any new type of LO (such as HDLO or HELO). As a matter of fact Energy Window Augmented Plane Wave (EWAPW) can be viewed as a competitor to the various forms of LO, as we argue that LO type wavefunctions will not significantly improve the accuracy of the basis set to obtain the eigenstates of the KS Hamiltonian (see Section III.6). Indeed, all of our basis, EWAPW, wavefunctions will be of the form $\frac{1}{\sqrt{\Omega}}\sum_{\mathbf{G}}\bar{o}_{\mathbf{k}\mathbf{G}}^{nj}\exp\left(i\left(\mathbf{k}+\mathbf{G}\right)\cdot\mathbf{r}\right)$ in the IR, for appropriate $\bar{o}_{\mathbf{k}\mathbf{G}}^{nj}$ as in $\left(LA\right)^{2}PW$. However, unlike $\left(LA\right)^{2}PW$ methods, inside the MT region we will also be modifying the linearization energy of the $j$’th basis element based on the eigenvalue information from the previous, $n-1$’st, iteration of the solution of the KS equations. More precisely $\bar{o}_{\mathbf{k}\mathbf{G}}^{nj}$ will be chosen so that in the IR the basis wavefunctions have the same plane wave expansion as eigenstates $\nu$ of the previous iterations with $\nu=j$. In the MT we will be augmenting these wavefunction at the energies $E_{j}^{n}$ which are nearly equal to the eigenenergies of the previous iteration $\varepsilon_{\mathbf{k}\nu}^{n-1}$ (with $j=\nu$) except we will be using a widowing function $\mathcal{E}_{w}^{n}\left(E\right)$ (which is a piecewise constant approximation to the identity function: $Id\left(E\right)=E$, see Fig. (1)) and linearizing (augmenting) at $E_{j}^{n}=\mathcal{E}_{w}^{n}\left(\varepsilon_{\mathbf{k}\nu}^{n-1}\right)\cong\varepsilon_{\mathbf{k}\nu}^{n-1}$. As such one needs solve the spherically averaged KS Hamiltonian at a discrete (on the order of ten to one hundred) number of values $\mathcal{E}_{w}^{n}\left(E\right)$ takes on - per angular momentum channel, making it practical. Furthermore as we are able to have many linearization energies the method is likely more accurate then APW+lo or LAPW (comparable to LAPW+LO+HDLO+HELO), however because we do not increase the basis size beyond APW levels its likely to have speeds comparable to APW+lo. Detailed derivations are given in Section II below.
@@
The rest of this paper is organized as follows. In Section II.1 we write down the EWAPW wavefunctions in several equivalent ways. This is the main result of this work. In Section II.2 we show how to incorporate EWAPW wavefunctions into a KS loop. In Section III we will discuss various technical details needed to make the procedure possible on modern computers. In Section IV we present technical steps to write down the KS Hamiltonian, overlap and electron density, this is highly similar to regular Full Potential Linearized Augmented Plane Waves (FLAPW). In Section V we conclude and various extensions are considered in the Appendices.
@@ II.1 Wavefunctions
Where $E_{1,n}^{U}=-\infty$, $E_{N,n}^{U}=+\infty$ and $E_{i-1,n}^{U}<E_{i,n}^{V}<E_{i,n}^{U}$ $\forall i$. Here $\Theta$ is the Heaviside function. Here $N-1$ , on the order of ten to one hundred, is the number of windows. The windowing function is a piecewise constant approximation to the identity function $Id\left(E\right)=E$. A typical $\mathcal{E}_{w}^{n}\left(E\right)$ is pictured in Fig. (1). We note that, as we shall see below, we will only need to solve Eq. (1) for the energies $E_{i,n}^{V}$. We further note that the windowing function divides the spectrum into non-overlapping intervals, windows, $E_{i-1,n}^{U}\leq E<E_{i,n}^{U}$ and assigns each window the value $E\rightarrow E_{i,n}^{V}$. Some methods to obtain $E_{i,n}^{V}$ and $E_{i,n}^{U}$ are described in Section (III.7). Here we simply note that, in general, it is advantageous to place many $E_{i,n}^{U}$ in the energy range of the valence and semicore bands (as such approximating the identity more accurately in those crucial regions) and a few for the rest of the spectrum (which is less crucial). We now introduce the wavefunctions: $\Phi_{\mathbf{k}\mathbf{G}}^{n}\left(E\right)=$
$\left\{\begin{array}[]{cc}\frac{1}{\sqrt{\Omega}}\exp\left(i\left(\mathbf{k}+\mathbf{G}\right)\cdot\mathbf{r}\right)&\mathbf{r}\in IR\\ \sum_{lm}A_{\mathbf{k}\mathbf{G}}^{lm\alpha n}\left(E\right)u_{l\alpha}^{n}\left(r,\mathcal{E}_{w}^{n}\left(E\right)\right)Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\alpha}}\right)&\mathbf{r}\in MT_{\alpha}\end{array}\right.$ (10)
@@
$A_{\mathbf{k}\mathbf{G}}^{lm\alpha n}\left(E\right)=\frac{\frac{1}{\sqrt{\Omega}}4\pi i^{l}j_{l}\left(\left|\mathbf{k}+\mathbf{G}\right|S_{\alpha}\right)Y_{lm}^{*}\left(\widehat{\mathbf{k}+\mathbf{G}}\right)}{u_{l\alpha}^{n}\left(S_{\alpha},\mathcal{E}_{w}^{n}\left(E\right)\right)}$ (11)
making the wavefunction continuous. We note that $u_{l\alpha}^{n}\left(r,\mathcal{E}_{w}\left(E\right)\right)=u_{l\alpha}^{n}\left(r,E_{i,n}^{V}\right)$ for some $E_{i,n}^{V}$ so that Eq. (1) need only be solved at a discrete (small) number of energy values to obtain the basis in Eq. (10). Now suppose that for the $n$th iteration of the solutions to the KS equations (see Eq. (17) below) we have obtained eigenstates of the KS equations, $\Psi_{\mathbf{k}\nu}^{n}$ with eigenvalues $\varepsilon_{\mathbf{k}\nu}^{n}$, of the form:
@@
for some energies $E^{n}\left(\nu,\mathbf{k},\mathbf{G}\right)$, whereby $\mathcal{E}_{w}^{n}\left(E^{n}\left(\nu,\mathbf{k},\mathbf{G}\right)\right)$ is in the set $\left\{E_{i,n}^{V}\right\}$. Then we that the EWAPW basis for the $n+1$’st iteration of the solution of the KS equations is given by:
@@
When $\mathcal{E}_{w}^{n+1}$ is nearly the identity function $\mathcal{E}_{w}^{n+1}\left(\varepsilon_{\mathbf{k}\nu}^{n}\right)\cong\varepsilon_{\mathbf{k}\nu}^{n}$ and the calculations are well converged $\varepsilon_{\mathbf{k}\nu}^{n+1}\cong\varepsilon_{\mathbf{k}\nu}^{n}$ this error is small, indeed there are multiple $E_{i,n}^{V}$ in the valence and semicore bands so the error $\left|\varepsilon_{\mathbf{k}\nu}^{n}-\mathcal{E}_{w}^{n+1}\left(\varepsilon_{\mathbf{k}\nu}^{n}\right)\right|$ is much smaller then bandwidth. We note that given that overall we have created a basis that is given by the previous eigenbasis in the interstitial and augmented it with the eigenstates of the spherically averaged KS Hamiltonian (with the linearization energy relevant of the energies of the previous basis, nearly $\varepsilon_{\mathbf{k}\nu}^{n}$ in the limit of many windows) with good convergence between iterations and good spherical symmetry in the MT spheres, in EWAPW, we are working with a basis of near eigenvectors. As such the KS Hamiltonian and overlap matrices will be nearly diagonal and sparse and as such quicker to diagonalize then the LAPW Hamiltonian and overlap matrices. We further note that unlike LAPW or APW+lo we have many radial wavefunctions $u_{l\alpha}^{n}\left(r,E_{i,n}^{V}\right)$ per angular momentum channel, it is a comparable or greater amount of radial wavefunctions to LAPW+LO+HDLO+HELO. Furthermore we note that we only have $N_{\mathbf{G}}$ basis wavefunctions per $\mathbf{k}$ point. Therefore it is reasonable to surmise that we have a basis with a comparable speed to LAPW+lo and comparable accuracy to LAPW+LO+HDLO+HELO.

### v6

@@ Abstract
In this work we present a new method for basis set generation for electronic structure calculations of crystalline solids. This procedure is aimed at applications to Density Functional Theory (DFT). In this construction, Energy Window Augmented Plane Waves (EWAPW), we take advantage of the fact that most DFT calculations use a convergence loop in order to obtain the self consistent eigenstates of the final (converged) Kohn Sham (KS) Hamiltonian. Here we propose that, for the basis used at each step of the self consistency iteration, we use the previous eigenstate basis, in the interstitial region, and augment it, inside each Muffin Tin (MT) sphere, with the solution to the spherically averaged KS Hamiltonian for the linearization energy of the energy window which contains the energy of that previous eigenstate. Indeed, to reduce the number of times the spherically averaged KS potential needs to be solved inside the MT spheres it is advantageous break up the spectrum into non-overlapping intervals, windows, and solve the spherically averaged KS Hamiltonian inside the MT region only once per window per angular momentum channel (at the linearization energy relevant to that window, usually near the middle of the window). For practical applications it is reasonable to have on the order of five to fifty windows. At each step of the iteration of the solution of the KS equations the EWAPW basis is that of near eigenstates of the KS Hamiltonian for that iteration. Overall the basis size is the comparable with the Augmented Plane Waves (APW) basis set but the number of radial wavefunctions is comparable or greater to Linearized Augmented Plane Waves + Local Orbitals + Higher Derivative Local Orbitals + High Energy Local Orbitals (LAPW+LO+HDLO+HELO).
@@ I Introduction
The usefulness of a basis set for all electron calculations of crystalline solids within Density Functional Theory (DFT) methods is mainly determined by 1) its linearization error (that is energy errors for eigenstates of the Kohn Sham (KS) Hamiltonian inside the valence band), 2) its ability to handle semi-core states - that is the situation when there is more then one relevant band (not just the valence band) just at or below the Fermi energy, 3) the effectiveness with which the basis set converges to the energy obtained in the infinite basis limit as a function of the number of basis elements for the solution of the KS equations, $N_{tot}$ or how hard or soft the basis is, 4) the efficiently with which one can write down the KS Hamiltonian in that basis, amongst other things [1, 2, 3]. There are now several known competitors, in the literature, for which basis set is optimal. These can be divided into mostly plane wave basis methods [1, 2, 8, 9, 5, 6, 3, 4, 7, 47] and mostly Muffin Tin (MT) basis sets [2, 11, 4, 12, 13, 15, 10, 14, 47]. In this work we focus on the plane wave methods [1, 2, 8, 9, 5, 6, 3, 4, 7]. Nominally (without augmentation or pseudization), within the plane wave method, the basis set is very simple but is extremely inefficient at obtaining good eigenstates of the KS Hamiltonian for reasonably small wave vector cutoffs - is extremely hard. Indeed near the nuclei, where the electron density is high and highly concentrated (oscillating), the KS potential has high wave vector components so that the eigenstates of the KS Hamiltonian have high wave vector components as well. This means that impractically large numbers of plane waves are needed to accurately describe the environment near atomic nuclei, making the situation generically numerically impossible. There are two main approaches to cure this problem - that is still use mainly plane wave basis sets: pseudopotentials [16, 17, 2, 3] and augmentation methods [4, 2, 18, 8, 9, 19, 20, 21, 7, 47]. In this work we focus on augmentation methods.
@@
An early solution to the glueing problem due to Slater was simply to chose the linearization in the middle of the valence band and augment with plane waves to obtain the following APW basis $\phi_{\mathbf{k}\mathbf{G}}^{n}\left(E^{n}\right)=$
$\left\{\begin{array}[]{cc}\frac{1}{\sqrt{\Omega}}\exp\left(i\left(\mathbf{k}+\mathbf{G}\right)\cdot\mathbf{r}\right)&\mathbf{r}\in IR\\ \sum_{lm}A_{\mathbf{k}\mathbf{G}}^{lm\alpha n}u_{l\alpha}^{n}\left(r,E^{n}\right)Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\alpha}}\right)&\mathbf{r}\in MT_{\alpha}\end{array}\right.$ (2)
Where $E^{n}$ is the $n$’th iteration linearization energy, $IR$ is the interstitial, $MT_{\alpha}$ is the $\alpha$th MT sphere, $\Omega$ is the volume of the unit cell, $\mathbf{k}$ is a point in the first Brillouin, $\mathbf{G}$ is a reciprocal lattice vector, $Y_{lm}$ are spherical harmonics, and $A_{\mathbf{k}\mathbf{G}}^{lm\alpha n}$ are the matching coefficients that insure the wavefunction is continuous. Furthermore the $A_{\mathbf{k}\mathbf{G}}^{lm\alpha n}$ can be computed using the relation [22]:
which decouples the matching problem into separate angular momentum channels. Here $j_{l}$ are the spherical Bessel functions. There is a cutoff $l_{max}\sim 8-12$, which comes about because of the negligibly small values of the spherical Bessel functions, for large $l$, inside the MT spheres. This is not an efficient solution because of the large linearization errors for the wavefunction. Alternatively it is possible to choose $E^{n}$ self consistently that is the same as every eigenvalue for each eigenwavefunction (however that is very numerically demanding) [2, 1].
One solution to this glueing problem is the Linearized Augmented Plane Waves (LAPW) basis set [4, 2, 1]. Here plane waves are augmented within the MT sphere by a linear combination of the solutions to the spherically averaged KS Hamiltonian (see Eq. (1)) at the linearization energy $E_{L}^{n}$ (typically the middle of the valence band) $u_{l\alpha}^{n}\left(r,E_{L}^{n}\right)$ and its derivative with respect to energy $\dot{u}_{l\alpha}^{n}\left(r,E_{L}^{n}\right)=\frac{\partial}{\partial E}u_{l\alpha}^{n}\left(r,E_{L}^{n}\right)$. This is done for all angular momentum channels $lm$ (up to $l_{max}$) and all atoms $\alpha$. Whereby, by matching the MT wavefunctions to the planewave wavefunction, that is demanding the total function be continuous and continuously differentiable - at the MT sphere radius [4, 2, 1, 5, 6] - we obtain a efficient smooth basis set. The LAPW basis set may be written as:
@@
Where $a_{\mathbf{k}\mathbf{G}}^{lm\alpha n}$ and $b_{\mathbf{k}\mathbf{G}}^{lm\alpha n}$ are the matching coefficients that insure the wavefunction is continuous and continuously differentiable. Furthermore $a_{\mathbf{k}\mathbf{G}}^{lm\alpha n}$ and $b_{\mathbf{k}\mathbf{G}}^{lm\alpha n}$ can be computed using the relation in Eq. (3). This allows for efficient description of the valence band which is adjustable to the form of the KS potential near the nuclei. This idea can be extended to Linearized Augmented Lattice Adapted Plane Waves ($\left(LA\right)^{2}PW$) which have the form [18]:
@@
for $\nu=j$. One then truncates the basis set with number of basis functions $j\ll N_{\mathbf{G}}$, as one has a basis of near eigenvectors in the IR, and runs the KS loop to convergence. This greatly reduces the basis set size needed for the iterative calculations after the first iteration [18]. These methods can be further improved as in Quadratically Augmented Plane Waves QAPW [20, 21] where $\ddot{u}_{l\mu}^{n}\left(r,E_{L}^{n}\right)=\frac{\partial^{2}}{\partial E^{2}}u_{l\mu}^{n}\left(r,E_{L}^{n}\right)$ is also used inside the MT sphere and one more derivative (with respect to the radial co-ordinate) is matched to the plane waves in the IR which reduces linearization errors. However, the more derivatives are matched the more the radial wavefunction inside the MT sphere look like Bessel functions (indeed in the limit where an infinite number of derivatives is matched the radial wavefunction must be exactly Bessel) the higher cutoff, the harder the basis becomes [7, 18]. Furthermore if there are several relevant bands one can use Local Orbitals (LO) basis wave functions [8, 1, 9, 23, 18, 7]. Many LO like basis wavefunctions are of the form:
$\phi_{LO}^{nlm\alpha}=\left\{\begin{array}[]{cc}0&\mathbf{r}\in IR,MT_{\beta}\>for\>\beta\neq\alpha\\ \left[\bar{a}^{l\alpha n}u_{l\alpha}^{n}\left(r,E_{L}^{n}\right)+\bar{b}^{l\alpha n}\dot{u}_{l\alpha}^{n}\left(r,E_{L}^{n}\right)+\bar{c}^{l\alpha n}\tilde{u}_{l\alpha}^{n}\left(r\right)\right]Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\alpha}}\right)&\mathbf{r}\in MT_{\alpha}\end{array}\right.$ (8)
Where the wavefunction is chosen to be continuous and continuously differentiable using the coefficients $\bar{a}^{l\alpha n}$, $\bar{b}^{l\alpha n}$ and $\bar{c}^{l\alpha n}$. There are many choices for $\tilde{u}_{l\alpha}^{n}$. For regular LO these are chosen as $\tilde{u}_{l\alpha}^{n}\left(r\right)=u_{l\alpha}^{n}\left(r,E_{SC}^{n}\right)$ where $SC$ stands for semicore states. For Higher Derivative Local Orbitals [23, 25, 7, 18, 24] (HDLO) $\tilde{u}_{l\alpha}^{n}=\ddot{u}_{l\alpha}^{n}\left(r,E_{L}^{n}\right)$, while for Higher Energy Local Orbitals [23, 25, 7, 18, 24] (HELO) the $\tilde{u}_{l\alpha}^{n}$ is an eigenstate of the spherically averaged KS Hamiltonian (inside $MT_{\alpha}$) with energy $E_{HELO}^{nl}$ such that
@@
A competitor to the LAPW method is the APW+lo method which uses the basis in Eq. (2) and augments it with lo wavefunctions of the form:
$\phi_{lo}^{nlm\alpha}=\left\{\begin{array}[]{cc}0&\mathbf{r}\in IR,MT_{\beta}\>for\>\beta\neq\alpha\\ \left[a_{lo}^{l\alpha n}u_{l\alpha}^{n}\left(r,E^{n}\right)+b_{lo}^{l\alpha n}\dot{u}_{l\alpha}^{n}\left(r,E^{n}\right)\right]Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\alpha}}\right)&\mathbf{r}\in MT_{\alpha}\end{array}\right.$
where the coefficients $a_{lo}^{l\alpha n}$ and $b_{lo}^{l\alpha n}$ are chosen to make the wavefunction continuous. These additional wavefunctions [9, 50, 46] which often come in a very small number of angular momentum channels [46, 9, 50] lead to a linearization of the APW basis set.
@@
In this work we will not be proposing any new type of LO (such as HDLO or HELO or conventional LO). As a matter of fact Energy Window Augmented Plane Wave (EWAPW) can be viewed as a competitor to the various forms of LO and lo, as we argue that LO and lo type wavefunctions will not significantly improve the accuracy of the basis set to obtain the eigenstates of the KS Hamiltonian (see Section III.6). Indeed, all of our basis, EWAPW, wavefunctions will be of the form $\frac{1}{\sqrt{\Omega}}\sum_{\mathbf{G}}\bar{o}_{\mathbf{k}\mathbf{G}}^{nj}\exp\left(i\left(\mathbf{k}+\mathbf{G}\right)\cdot\mathbf{r}\right)$ in the IR, for appropriate $\bar{o}_{\mathbf{k}\mathbf{G}}^{nj}$ as in $\left(LA\right)^{2}PW$. However, unlike $\left(LA\right)^{2}PW$ methods, inside the MT region we will also be modifying the linearization energy of the $j$’th basis element based on the eigenvalue information from the previous, $n-1$’st, iteration of the solution of the KS equations. More precisely $\bar{o}_{\mathbf{k}\mathbf{G}}^{nj}$ will be chosen so that in the IR the basis wavefunctions have the same plane wave expansion as eigenstates $\nu$ of the previous iterations with $\nu=j$. In the MT we will be augmenting these wavefunction at the energies $E_{j}^{n}$ which are nearly equal to the eigenenergies of the previous iteration $\varepsilon_{\mathbf{k}\nu}^{n-1}$ (with $j=\nu$) except we will be using a widowing function $\mathcal{E}_{w}^{n}\left(E\right)$ (which is a piecewise constant approximation to the identity function: $Id\left(E\right)=E$, see Fig. (1)) and linearizing (augmenting) at $E_{j}^{n}=\mathcal{E}_{w}^{n}\left(\varepsilon_{\mathbf{k}\nu}^{n-1}\right)\cong\varepsilon_{\mathbf{k}\nu}^{n-1}$. As such one needs solve the spherically averaged KS Hamiltonian at a discrete (on the order of five to fifty) number of values $\mathcal{E}_{w}^{n}\left(E\right)$ takes on - per angular momentum channel, making it practical. Furthermore as we are able to have many linearization energies the method is likely more accurate then APW+lo or LAPW (comparable to LAPW+LO+HDLO+HELO), however we do not increase the basis size beyond APW levels. Detailed derivations are given in Section II below. We now note that the numerical efficiency of a DFT calculation is often determined by the diagonalization time and the time it takes to numerically set up the KS Hamiltonian and charge density [34]. Overall EWAPW has roughly the same diagonalization time as APW (same size of basis) but takes a comparable amount of time to set up the charge density and KS Hamiltonian as LAPW+LO+HDLO+HELO as it has a comparable number of radial basis functions. We note that in many situations with a large basis set the diagonalization time is the biggest [34].
@@
The rest of this paper is organized as follows. In Section II.1 we write down the EWAPW wavefunctions in several equivalent ways. This is the main result of this work. In Section II.2 we show how to incorporate EWAPW wavefunctions into a KS loop. In Section III we will discuss various technical details needed to make the procedure possible on modern computers as well as some technical comments useful for computational efficiency. In Section IV we present technical steps to write down the KS Hamiltonian, overlap and electron density, this is highly similar to regular Full Potential Linearized Augmented Plane Waves (FLAPW). In Section V we conclude. Various extensions are considered in the Appendices.
@@ II.1 Wavefunctions
Where $E_{1,n}^{U}=-\infty$, $E_{N,n}^{U}=+\infty$ and $E_{i-1,n}^{U}<E_{i,n}^{V}<E_{i,n}^{U}$ $\forall i$. Here $\Theta$ is the Heaviside function. Here $N-1$ , on the order of five to fifty, is the number of windows. The windowing function is a piecewise constant approximation to the identity function $Id\left(E\right)=E$. A typical $\mathcal{E}_{w}^{n}\left(E\right)$ is pictured in Fig. (1). We note that, as we shall see below, we will only need to solve Eq. (1) for the energies $E_{i,n}^{V}$. We further note that the windowing function divides the spectrum into non-overlapping intervals, windows, $E_{i-1,n}^{U}\leq E<E_{i,n}^{U}$ and assigns each window the value $E\rightarrow E_{i,n}^{V}$. Some methods to obtain $E_{i,n}^{V}$ and $E_{i,n}^{U}$ are described in Section (III.7). Here we simply note that, in general, it is advantageous to place many $E_{i,n}^{U}$ in the energy range of the valence and semicore bands (as such approximating the identity more accurately in those crucial regions) and a few for the rest of the spectrum (which is less crucial). We now introduce the wavefunctions: $\Phi_{\mathbf{k}\mathbf{G}}^{n}\left(E\right)=$
$\left\{\begin{array}[]{cc}\frac{1}{\sqrt{\Omega}}\exp\left(i\left(\mathbf{k}+\mathbf{G}\right)\cdot\mathbf{r}\right)&\mathbf{r}\in IR\\ \sum_{lm}\bar{A}_{\mathbf{k}\mathbf{G}}^{lm\alpha n}\left(E\right)u_{l\alpha}^{n}\left(r,\mathcal{E}_{w}^{n}\left(E\right)\right)Y_{lm}\left(\widehat{\mathbf{r}-\mathbf{r}_{\alpha}}\right)&\mathbf{r}\in MT_{\alpha}\end{array}\right.$ (11)
@@
$\bar{A}_{\mathbf{k}\mathbf{G}}^{lm\alpha n}\left(E\right)=\frac{\frac{1}{\sqrt{\Omega}}4\pi i^{l}j_{l}\left(\left|\mathbf{k}+\mathbf{G}\right|S_{\alpha}\right)Y_{lm}^{*}\left(\widehat{\mathbf{k}+\mathbf{G}}\right)}{u_{l\alpha}^{n}\left(S_{\alpha},\mathcal{E}_{w}^{n}\left(E\right)\right)}$ (12)
making the wavefunction continuous. We note that:
$\Phi_{\mathbf{k}\mathbf{G}}^{n}\left(E\right)=\phi_{\mathbf{k}\mathbf{G}}^{n}\left(\mathcal{E}_{w}^{n}\left(E\right)\right)$ (13)
We note that $u_{l\alpha}^{n}\left(r,\mathcal{E}_{w}\left(E\right)\right)=u_{l\alpha}^{n}\left(r,E_{i,n}^{V}\right)$ for some $E_{i,n}^{V}$, so that Eq. (1) need only be solved at a discrete (small) number of energy values to obtain the basis in Eq. (11). Now suppose that for the $n$th iteration of the solutions to the KS equations (see Eq. (19) below) we have obtained eigenstates of the KS equations, $\Psi_{\mathbf{k}\nu}^{n}$ with eigenvalues $\varepsilon_{\mathbf{k}\nu}^{n}$, of the form:
@@
for some energies $E^{n}\left(\nu,\mathbf{k},\mathbf{G}\right)$. Then we that the EWAPW basis for the $n+1$’st iteration of the solution of the KS equations is given by:
@@
When $\mathcal{E}_{w}^{n+1}$ is nearly the identity function $\mathcal{E}_{w}^{n+1}\left(\varepsilon_{\mathbf{k}\nu}^{n}\right)\cong\varepsilon_{\mathbf{k}\nu}^{n}$ and the calculations are well converged $\varepsilon_{\mathbf{k}\nu}^{n+1}\cong\varepsilon_{\mathbf{k}\nu}^{n}$ this error is small, indeed there are multiple $E_{i,n}^{V}$ in the valence and semicore bands so the error $\left|\varepsilon_{\mathbf{k}\nu}^{n}-\mathcal{E}_{w}^{n+1}\left(\varepsilon_{\mathbf{k}\nu}^{n}\right)\right|$ is much smaller then bandwidth. We note that given that overall we have created a basis that is given by the previous eigenbasis in the interstitial and augmented it with the eigenstates of the spherically averaged KS Hamiltonian (with the linearization energy relevant of the energies of the previous basis, nearly $\varepsilon_{\mathbf{k}\nu}^{n}$ in the limit of many windows) with good convergence between iterations and good spherical symmetry in the MT spheres, in EWAPW, we are working with a basis of near eigenvectors. As such the KS Hamiltonian and overlap matrices will be nearly diagonal and sparse and as such quicker to diagonalize then the LAPW Hamiltonian and overlap matrices. We further note that unlike LAPW or APW+lo we have many radial wavefunctions $u_{l\alpha}^{n}\left(r,E_{i,n}^{V}\right)$ per angular momentum channel, it is a comparable or greater amount of radial wavefunctions to LAPW+LO+HDLO+HELO. Furthermore we note that we only have $N_{\mathbf{G}}$ basis wavefunctions per $\mathbf{k}$ point. Therefore it is reasonable to surmise that we have a basis with a comparable diagonalization speed to APW and comparable accuracy to LAPW+LO+HDLO+HELO. We note that setting up the KS Hamiltonian, overlap and charge density will take comparable time to LAPW+LO+HDLO+HELO methods.